You are planning to purchase a car in three years and wish to have $5,000 for a down payment. You have access to an account that earns 10%, compounded monthly. How much do you need to deposit into the account each month for the next three years in order to reach your goal? $161.75 $119.67 $3,756.57 $3,708.70
Answer #1

Interest Rate =10 % compounded monthly, Applicable Monthly Rate = 10/12 =0.8333 %, Tenure = 3 years or (3 x 12) = 36 months, Target Future Value = $ 5000

Let the monthly deposits be $ K

Therefore, 5000 = K x (1.008333)^(35) + ...........+ K x (1.008333)^(2) + K x (1.00833) + K = K x [{(1.008333)^(36)-1}/{1.008333-1}] = K x 41.782

K = 5000 / 41.782 = $ 119.67

Hence, the correct option is (b)